OpenStudy (anonymous):

Simplify: (x^2 + 2x +1)/(2x + 2)....Plz HELP

OpenStudy (anonymous):

\[\frac{x^2+2x+1}{2x+2}\] \[\frac{(x+1)(x+1)}{2(x+1)}\] \[\frac{(x+1)}{2}\] Factor the top and you'll be able to cancel x+1 from top and bottom.

OpenStudy (anonymous):

(x+1)/2 is right

OpenStudy (anonymous):

Okay i see, thanks.